# Flight Demonstration and Model Validation of a Prototype Variable-Altitude Venus Aerobot

### v1

@@ 5 Discussion: Application to Venus
The primary utility of FLOATS after validation against Earth flight test data is to predict flight trajectories on Venus, allowing simulation of a full-scale aerobot to aid mission planning studies. On Venus, the altitude control capability would be used by the science operations team to interrogate multiple altitudes over different times of day, allowing for a extensive characterization of the cloud layer with the onboard instrumentation. For supporting Venus simulations, the Venus Climate Database (VCD) [32] has been integrated into FLOATS. The VCD is a comprehensive compilation of atmospheric data and models for Venus, encompassing information on the planet’s atmospheric composition, structure, and dynamics, as well as data on surface features and climate patterns. Solar and infrared fluxes are taken from Robinson and Crisp [33] as a function of local solar zenith angle.
@@
Current estimates from envelope coupon helium retention measurements are consistent with a 1-Venus-day flight duration (117 Earth days), but minimum science baseline would be closer to a few weeks. One such baseline Venus flight of a 12.5 meter diameter aerobot (taken from Hall et al. [5] case B2) is simulated in Figure 15, illustrating three circumnavigations of the planet.
@@
Figure 15: Simulated trajectory for a 12 meter diameter aerobot on Venus (100 kg payload capacity) for 25 Earth days. Left: Aerobot latitude and longitude assuming zonal and meridional winds from VCD and a sunset entry at $5^{\circ}$ latitude. Right: Altitude over the Venus surface given a venting action after the first circumnavigation and a pumping action after the third local noon.
First, the passive altitude dynamics of the aerobot for global scale Venus winds and fluxes appear stable, especially at night. Solar heating during the day however does create an excess in ZP buoyancy, as the balloon material still retains a non-negligible $\alpha/\epsilon$ ratio. The equilibrium altitude accordingly tracks a kilometer-scale excursion as a function of the solar zenith angle peaking at local noon, which could be mitigated if desired through closed-loop altitude control.
@@ 6 Conclusion
    Prototype Fabrication: We describe our aerobot which is fabricated from Venus-relevant materials — building upon smaller prototypes in prior work [13, 5] and showing an increase in balloon scale and manufacturing fidelity. The aerobot in this flight test includes a metalized bilaminate of Teflon® and Kapton® for the outer ZP envelope, and a Vectran® woven inner SP envelope with urethane bladder. These materials are the flight materials expected for surviving Venus thermal conditions of $60^{\circ}$ [6], >94% concentration sulfuric acid aerosols [12], and the intense solar heating of 2300 W/m2 [33].
@@
    Buoyancy Modulation through Helium Gas Transfer: We demonstrate the ability to move helium gas with a pumping/venting system between balloon chambers to execute these altitude control maneuvers. This adapts the balloon-in-balloon architecture described in Voss [14] to a Venus design, protecting the high-pressure envelope from the acid aerosols in the external atmosphere. The buoyancy modulation authority is shown to be sufficient to vary altitude in the aerobot free-flight, and is further validated by stationary measurements given in the supplemental material.
@@
    Flight Instrumentation: We describe and implement an instrumentation set for capturing the aerobot dynamics and thermodynamics in sufficient detail for deriving a first-principal dynamic model of the aerobot. This includes barometers, thermistors, radiometers, cameras, and GPS trackers; as well as piloting infrastructure from the command center. While this instrumentation does not yet utilize spaceflight-grade components, it provides a template for the Venus avionics of future work.
@@
Appendix
@@ 6.1 Avionics and Sensor Calibration
Figure 16: Avionics diagram of the instrumentation and gas transfer system split over four platforms. Arrows show direction of information flow between different components with arrow labels indicating protocol for information transfer or the nature of the connection. Images below the avionics diagram show the common instrumentation stack with the pumping and power board and the instrument interface board.
@@ 6.2 Instrument Calibration
We performed post-flight calibration of several instruments and incorporated any corrections and biases resulting from the calibration into data recorded during flight. Figures 17-19 show results from the calibration of the SP transducer, thermistors, and anemometer. The Paroscientific microbarometer and the radiation sensors are supplied pre-calibrated from manufacturers along with a calibration sheet. The SP transducer was calibrated against a NIST-standard barometer by varying the test pressure across the full dynamic range of the SP transducer in the ascending and descending directions. Thermistors were thermally coupled with a large copper plate, with the cables in the flight configuration. The temperature of the copper plate was assumed to be uniform and the relative difference between the thermistors was noted. Air temperature was also recorded using the ultrasonic anemometer. For anemometer calibrations, the instrument was placed indoors in a windless room. After allowing transients to settle (approximately 90 minutes), the values of wind velocity along the three axes were noted.
@@
Figure 18: Calibration of the thermistors. All thermistors were attached flush with a copper plate using copper tape to ensure good thermal contact. Each thermistor had the same length of cable as in the flights. Measurements show the presence of relative bias in the thermistors – these result from manufacturing uncertainties and also capacitance from the length of cable attached to each thermistor, which could be longer than 10 meters for the top thermistor.
@@ 6.3 Buoyancy force that contributes to initial balloon film stress
Figure 20: Derivation Diagram for Base Section Buoyancy: The SP weight, minus it’s own buoyancy within the ZP balloon, can be imagined to be supported by a support truss structure that creates the bottom spherical shape. The buoyancy of that shape is the given by a surface integral pressure difference
@@
Let’s use spherical coordinates to parameterize the balloon,
@@
where $\psi\in[0,2\pi)$ is an angle about the vertical axis and $\phi\in[0,\pi]$ is from the SP apex. In our case, the axisymmetric assumption means that the forces in the $x$ and $y$ directions will be zero, so all we need to worry about are the forces in the $z$ direction.
@@ Acknowledgments
The research described in this paper was funded by the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration. The authors would like to thank: Kirk Barrow, Charlene Paloma, Tanya Samra, Faustino Chirino, and Lance Christensen for making our field tests a reality; Ahyde Lara, Rich Bauer, and Allen Dial for supporting the aerobot fabrication and flight; the Black Rock Nevada BLM Office and the Tillamook Air Museum for the use of their property; Satish Khanna, Chris Yahnker, Pat Beauchamp, Kim Aaron, Peter Dillon, Dara Sabahi, Michael Meacham, John Gallon, Katie Siegel, Michael Schein, and Jim Denman for their technical guidance of the program; Bill Warner, Dominic Aldi, Jared Leblanc, and Emma Bradford for the material testing of balloon envelopes; and Blair Emanuel and Tonya Beatty for the post-flight evaluation of the aerobot prototype after the Nevada field campaign.

### v2

@@ 5 Discussion: Application to Venus
The primary utility of FLOATS after validation against Earth flight test data is to predict flight trajectories on Venus, allowing simulation of a full-scale aerobot to aid mission planning studies. On Venus, the altitude control capability would be used by the science operations team to interrogate multiple altitudes over different times of day, allowing for a extensive characterization of the cloud layer with the onboard instrumentation. For supporting Venus simulations, the Venus Climate Database (VCD) [34] has been integrated into FLOATS. The VCD is a comprehensive compilation of atmospheric data and models for Venus, encompassing information on the planet’s atmospheric composition, structure, and dynamics, as well as data on surface features and climate patterns. Solar and infrared fluxes are taken from Robinson and Crisp [35] as a function of local solar zenith angle and applied purely as diffuse sources, as little direct sunlight reaches the target altitudes.
@@
Current estimates from envelope coupon helium retention measurements are consistent with a 1-Venus-day flight duration (117 Earth days), but minimum science baseline would be closer to a few weeks. One such baseline Venus flight of a 12.5 meter diameter aerobot (taken from Hall et al. [5] case B2) is simulated in Figure 15, illustrating three circumnavigations of the planet after entry at $5^{\circ}$ latitude at local sunset with zonal and meridional winds imported from VCD.
@@
Figure 15: Simulated trajectory for a 12 meter diameter aerobot on Venus (100 kg payload capacity) for 25 Earth days. Left: Aerobot latitude and longitude. Right: Altitude over the Venus surface given a venting action after the first circumnavigation and a pumping action after the third local noon.
First, the passive altitude dynamics of the aerobot for global scale Venus winds and fluxes appear stable, especially at night. Solar heating during the day however does create an excess in ZP buoyancy, as the balloon material still retains a non-negligible $\alpha/\epsilon$ ratio. The equilibrium altitude accordingly tracks a kilometer-scale excursion as a function of the solar zenith angle peaking at local noon, which could be mitigated if desired through closed-loop altitude control in the BCM.
@@ 6 Conclusion
    Prototype Fabrication: We describe our aerobot which is fabricated from Venus-relevant materials — building upon smaller prototypes in prior work [13, 5] and showing an increase in balloon scale and manufacturing fidelity. The aerobot in this flight test includes a metalized bilaminate of Teflon® and Kapton® for the outer ZP envelope, and a Vectran® woven inner SP envelope with urethane bladder. These materials are the flight materials expected for surviving Venus thermal conditions of $60^{\circ}$C [7], >94% concentration sulfuric acid aerosols [12], and the intense solar heating of 2300 W/m2 [35].
@@
    Buoyancy Modulation through Helium Gas Transfer: We demonstrate the ability to move helium gas with a pumping/venting system between balloon chambers to execute these altitude control maneuvers. This adapts the balloon-in-balloon architecture described in Voss [14] to a Venus design, protecting the high-pressure envelope from the acid aerosols in the external atmosphere. The buoyancy modulation authority is shown to be sufficient to vary altitude in the aerobot free-flight.
@@
    Flight Instrumentation: We describe and implement an instrumentation set for capturing the aerobot dynamics and thermodynamics in sufficient detail for deriving a first-principal dynamic model of the aerobot. This includes barometers, temperature sensors, radiometers, cameras, and GPS trackers; as well as piloting infrastructure from the command center. While this instrumentation does not yet utilize spaceflight-grade components, it provides a template for the Venus avionics of future work.
@@
Appendix: Sensor Details and Shape Model Supplement
@@ 6.1 Avionics and Sensor Calibration
Figure 16: Avionics diagram of the instrumentation and gas transfer system split over four platforms. Images show the common instrumentation stack with the pumping and power board and the instrument interface board.
@@ 6.2 Instrument Calibration
We performed post-flight calibration of several instruments and incorporated any corrections and biases resulting from the calibration into data recorded during flight. Figures 17-19 show results from the calibration of the SP transducer, temperature sensors, and anemometer. The Paroscientific microbarometer and the radiation sensors are supplied pre-calibrated from manufacturers along with a calibration sheet. The SP transducer was calibrated against a NIST-standard barometer by varying the test pressure across the full dynamic range of the SP transducer in the ascending and descending directions. Temperature sensors were thermally coupled with a large copper plate, with the cables in the flight configuration. The temperature of the copper plate was assumed to be uniform and the relative difference between the temperature sensors was noted. Air temperature was also recorded using the ultrasonic anemometer. For anemometer calibrations, the instrument was placed indoors in a windless room. After allowing transients to settle (approximately 90 minutes), the values of wind velocity along the three axes were noted.
@@
Figure 18: Calibration of the temperature sensors. All sensors were attached to a copper plate using copper tape to ensure good thermal contact. Each sensor had the same length of cable as in the flights. Measurements show relative bias from manufacturing uncertainties and capacitance from the length of cable.
@@ 6.3 Buoyancy force that contributes to initial balloon film stress
Figure 20: Derivation Diagram for Base Section Buoyancy: The SP weight, minus it’s own buoyancy within the ZP balloon, can be imagined to be supported by a support truss structure that creates the bottom spherical shape. The buoyancy of that shape is the given by a surface integral pressure difference.
@@
Let’s use spherical coordinates to parameterize the contact area of the ZP and SP balloon envelopes, centered at the SP center,
@@
where $\psi\in[0,2\pi)$ is an angle about the vertical axis and $\phi\in[0,\pi]$ is the angle from the SP apex. In our case, the axisymmetric assumption means that the forces in the $x$ and $y$ directions will be zero, so all we need to worry about are the forces in the $z$ direction.
@@ Acknowledgments
The research described in this paper was funded by the Jet Propulsion Laboratory, California Institute of Technology, under contract 80NM0018D0004 with the National Aeronautics and Space Administration. The authors would like to thank: Kirk Barrow, Charlene Paloma, Tanya Samra, Faustino Chirino, and Lance Christensen for making our field tests a reality; Ahyde Lara, Richard Bauer, and Allen Dial for supporting the aerobot fabrication and flight; the Black Rock Nevada BLM Office and the Tillamook Air Museum for the use of their property; Satish Khanna, Chris Yahnker, Pat Beauchamp, Kim Aaron, Peter Dillon, Dara Sabahi, Michael Meacham, John Gallon, Katie Siegel, Michael Schein, and Jim Denman for their technical guidance of the program; Bill Warner, Dominic Aldi, Jared Leblanc, and Emma Bradford for the material testing of balloon envelopes; Spencer Diehl for reprocessing FLOATS simulations after first review, and Blair Emanuel and Tonya Beatty for the post-flight evaluation of the aerobot prototype after the Nevada field campaign.